OpenStudy (anonymous):

1 x 4x-9 _ - _ = ______ 4 9 9 Solve for x

OpenStudy (anonymous):

the GCD is 36. So, multiply both sides of the equation by 36 to get \[9 - 4x = 4(4x-9)\] simplify the right hand side to get \[9-4x = 16x -36\] add a 4x to both sides to get \[9 = 20x -36\] now add a 36 to both sides to get \[45 = 20x \] divide both sides by 20 to get the value of x